OpenStudy (anonymous):

The new business building uptown is building a wall made entirely of glass pyramids. Each pyramid will have a square base with length of 8 inches and height of 10 inches. If the wall requires 250 of these pyramids, what volume of glass will be used? Round to the nearest cubic inch.

OpenStudy (anonymous):

If a pyramid has a base of B square units, and a height of h units, then the a Pyramid volume V is 1/3 B x h cubic units, or V = 1/3 Bh ... the Base has to be AREA so with a square base you would need to square the side length to be in squared form for this it needs to be V=1/3(8)^2 X (10) = 1/3(64)(10) ............ = 21.12 x 10 = ..........211cm^3

OpenStudy (anonymous):

v=a^2*h/3 so v=8^2*10/3

OpenStudy (anonymous):

do 10/3 first @Aliypop

OpenStudy (anonymous):

3.33333333

OpenStudy (anonymous):

yep now do 8*8

OpenStudy (anonymous):

64

OpenStudy (anonymous):

now do 64*3.333333333333333333333333333333333333333

OpenStudy (anonymous):

213.333333

OpenStudy (anonymous):

yep do 213.333333*250

OpenStudy (anonymous):

53,333.33325

OpenStudy (anonymous):

yes which rounds to 53,333